# GAF: Gaussian Avatar Reconstruction from Monocular Videos via Multi-view Diffusion

### v1

@@ Abstract
We propose a novel approach for reconstructing animatable 3D Gaussian avatars from monocular videos captured by commodity devices like smartphones. Photorealistic 3D head avatar reconstruction from such recordings is challenging due to limited observations, which leaves unobserved regions under-constrained and can lead to artifacts in novel views. To address this problem, we introduce a multi-view head diffusion model, leveraging its priors to fill in missing regions and ensure view consistency in Gaussian splatting renderings. To enable precise viewpoint control, we use normal maps rendered from FLAME-based head reconstruction, which provides pixel-aligned inductive biases. We also condition the diffusion model on VAE features extracted from the input image to preserve details of facial identity and appearance. For Gaussian avatar reconstruction, we distill multi-view diffusion priors by using iteratively denoised images as pseudo-ground truths, effectively mitigating over-saturation issues. To further improve photorealism, we apply latent upsampling to refine the denoised latent before decoding it into an image. We evaluate our method on the NeRSemble dataset, showing that GAF outperforms the previous state-of-the-art methods in novel view synthesis by a 5.34% higher SSIM score. Furthermore, we demonstrate higher-fidelity avatar reconstructions from monocular videos captured on commodity devices. Project Page: https://tangjiapeng.github.io/projects/GAF
@@ 1 Introduction
In order to exploit multi-view diffusion priors for Gaussian head reconstruction, we employ a diffusion loss that uses iteratively denoised images as pseudo-ground truths for novel view renderings, instead of using a single-step score distillation sampling loss [45]. Moreover, to improve the fidelity of the Gaussian renderings, we introduce a latent upsampler model to enhance facial details within the denoised latent before decoding it back to image space. As demonstrated in Fig. 1, our approach generates high-quality, photorealistic novel views of head avatars using only short monocular videos captured by a smartphone. Extensive comparisons with state-of-the-art methods show that our approach delivers higher fidelity and more view-consistent avatar reconstructions from monocular videos. Our GAF significantly outperforms state-of-the-art methods, achieving 5.34% higher SSIM in novel view synthesis on the NeRSemble dataset.
Our contributions can be summarized as follows:
@@
  • We present a mechanism for enhancing the photorealism and cross-view consistency of Gaussian splatting by integrating latent upsampling and multi-view diffusion priors.
@@ 2 Related Work
To create photo-realistic animatable head avatars, several works [14, 91, 75, 57, 48, 76] have combined NeRF/GS with 3D morphable models (3DMM). NerFace [14] learned expression-dependent NeRF using head parameters from monocular videos. INSTA [91] and H3QAvatar [66] mapped query points to a canonical space via deformation fields and defined multi-resolution hash grids for head radiance fields. PointAvatar [88] explored point-based representations for efficient training. Our work is closely related to animatable Gaussian splats [48, 57], which attached splats to the triangles of a FLAME mesh, and updated their properties by triangle deformations controlled by FLAME parameters. Although promising results have been achieved, they typically require multi-view setups with high-quality cameras in studio environments [36, 27]. Some works reconstruct avatars from monocular videos [88, 1, 91, 75]. However, they only reconstruct the visible region from inputs, lacking effective priors to complete missing areas. To address this, we utilize multi-view diffusion priors to complete unobserved regions of the face, ensuring photorealistic renderings from extremely hold-out viewpoints.
@@
Researchers have demonstrated that large-scale pretrained text-to-image priors [52, 55, 51] can be distilled for 3D asset generation [31, 63] using the score distillation sampling (SDS) loss proposed by DreamFusion [45] and its variants [72, 81, 39]. Several studies [20, 6, 90] applied SDS loss for text-driven 3D head avatar generation. For single-view 3D reconstruction, approaches like RelFusion [39], Magic123 [46], and Dreambooth3D [50] adapt text-to-image diffusion priors to specific objects to preserve identity. However, achieving consistent novel views from text prompts alone remains challenging without leveraging the input image. To address this, we choose to learn multi-view image diffusion priors conditioned on a single image. Rather than using a single-step SDS loss, we use iteratively denoised images as pseudo-ground truths for novel view supervision, similar to ReconFusion [74]. By integrating latent upsampler diffusion with joint multi-view latent denoising, we enhance both multi-view consistency and fidelity.
@@
Instead of text-to-image diffusion priors, some works learn image-conditioned novel view diffusion priors, which leverage the identity and appearance details of the input image to generate consistent novel views. 3DiM [73] pioneered a diffusion model for pose-conditioned novel view generation. Zero-123 finetune StableDiffusion [52] on Objaverse [10, 11] dataset to improve generalization ability. Some works learn to generate consistent multiple views from a single image via introducing dense 3D self-attention or epipolar attention in the denoiser network, including MVDiffusion [65], SyncDreamer [35], Zero-123++ [58], Wonder3D [37], MVDream [59], ImageDream [71], Human3Diffusion [79], IM3D [40], CAT3D [15], and VideoMV [92]. These works primarily target general objects, whereas our focus is specifically on human heads. To obtain more specialized priors, we train multi-view diffusion models on a dedicated multi-view head video dataset. By using normal maps from FLAME mesh reconstruction as the camera condition, we achieve precise viewpoint control for head rendering. This pixel-aligned inductive bias introduced by normal maps enables more accurate novel view synthesis, crucial for consistent head Gaussian supervision.
@@ 3.2 Gaussian Avatars
$\vec{r}^{\prime}=\vec{R}\vec{r}$ (1)
$\vec{\mu}^{\prime}=k\vec{R}\vec{\mu}+\vec{T}$ (2)
$\vec{s}^{\prime}=k\vec{s}$ (3)
To capture fine-grained details in the rendered output, an adaptive density control strategy dynamically creates new Gaussians for each triangle based on screen-space positional gradients. This densification process operates within local space, where newly added Gaussians inherit binding relations from their original counterparts.
@@ 4.1 Normal map-conditioned Multi-view Head Latent Diffusion
To address missing regions in monocular videos with limited head coverage, one solution is to distill pre-trained text-to-image diffusion models to regularize novel view renderings. Personalized techniques like Dreambooth [54] allow for identity preservation by customizing diffusion models for specific objects we wish to reconstruct. However, since monocular videos may lack comprehensive side-view coverage, these models are prone to view bias: generated images tend to default to front-view perspectives, making it challenging to produce plausible novel views. Nevertheless, personalized text-to-image diffusion models are not designed to capture the distribution of novel views from a single input image. To overcome this, we propose a novel view diffusion model that generates identity-preserved and appearance-coherent novel views conditioned on an input image. By denoising multiple novel views simultaneously, our approach enhances cross-view consistency. The multi-view head diffusion model is illustrated in Fig. 3.
@@
Figure 3: Multi-view latent head diffusion models. Given multi-view noisy image latents, we concatenate them with VAE latents of normal maps rendered from FLAME tracking. These combined inputs are processed by a 2D U-Net denoiser with attention blocks. To maintain 3D consistency, 3D attention blocks apply cross-attention across all views, integrating face identity and appearance details from the input image into the denoised latents while allowing information exchange between noisy latents across views. Additionally, the CLIP embedding of the input image is used for global control via cross-attention blocks.
@@ 4.2 Gaussian Avatars Reconstruction with Multi-view Diffusion Priors
We now seek to utilize the multi-view diffusion priors for head Gaussian reconstruction. A commonly used strategy is Score Distillation Sampling (SDS) loss [45]. Let us denote the rendered image from Gaussian as $\vec{x}$. One can add noise $\vec{\epsilon}$ at time step $t$ to the input image $\vec{x}$ to obtain a noisy image $\vec{x}_{t}$, which is fed into denoising network to estimate out the noise $\vec{\hat{\epsilon}}$. The SDS loss calculates the difference $\|\vec{\hat{\epsilon}}-\vec{\epsilon}\|_{2}^{2}$. The gradients can be back-propagated to update 3D representation. The SDS loss essentially performs one-step denoising. Due to the stochastic nature of the denoising process introduced by random noise levels and seeds, it contains noisy gradients that disturb 3D optimization. Consequently, it often causes over-saturated appearance issues in synthesized 3D assets. Wu et al. [74] found that after iteratively denoising $\vec{x}_{t}$ for multiple steps, we can obtain a deterministic output $\vec{x}_{0}$. Based on this observation, we can calculate $\|\vec{x}_{0}$ $\vec{x}\|_{2}^{2}$ as the diffusion loss for 3D Gaussian optimization.
@@
At each iteration, we randomly select the $i$-th input frame $\vec{I}_{i}$ and its FLAME mesh $\mathcal{M}_{i}$. By randomly sampling 4 viewpoints $\{\phi_{j}\}_{j=1}^{4}$, then we can generate 4 random novel views from $\vec{I}_{i}^{\operatorname{view}}=\mathcal{R}(\mathcal{O}_{i},\{\phi_{j}\})$, and render normal maps $\mathbf{N}_{i}=\mathcal{R}(\mathcal{M},\{\phi_{j}\})$ from $\mathcal{M}_{i}$. We can employ normal map-guided multi-view diffusion priors to regularize $\vec{I}_{i}^{\operatorname{view}}$. Concretely, we encode $\vec{I}_{i}^{\operatorname{view}}$ into latent features $\vec{z}$, which is perturbed with noise $\epsilon_{t}$ to obtain noisy feature $\vec{z}_{t}$. The noise intensity of $\epsilon_{t}$ is controlled by the diffusion time step $t\sim[0.02,0.98]$. We then iteratively perform multiple denoising steps of latent diffusion until the final clean latent $\vec{z}_{0}$ is obtained. To accelerate the generation speed, we adopt the DDIM sampling strategy [60] by running $t/k$ intermediate steps, which can reduce the denoising steps by $k=20$ times. Then $\vec{z}_{0}$ is decoded back to 4 images $\vec{\hat{I}}_{i}^{j}$ that can be served as pseudo supervision.
@@
To further enhance the facial details in $\vec{\hat{I}}_{i}^{\operatorname{view}}$, we use a latent upsampler diffusion model to super-resolve the denoised 4-view latents $\vec{z}_{0}$, from a resolution of 32 $\times$ 32 to 64 $\times$ 64. This super-resolution step allows the pseudo-image ground truths $\vec{\hat{I}}_{i}^{j}$ to reach a final resolution of 512$\times$512. We use 10-step DDIM sampling in our latent upsampler inference.
@@ 4.3 Loss Functions
$\vspace{-1.5mm}\mathcal{L}_{\operatorname{img}}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{SSIM}+\lambda_{3}\mathcal{L}_{LPIPS}$ (6)
where $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.1$. We also introduce splat position and scale regularization terms to penalize abnormally distributed splats. The position regularization term ensures that Gaussians remain close to their attached triangles during optimization through:
It will be disabled when the local scale of the Gaussian w.r.t the attached triangle is less than $\epsilon_{pos}=0.6$. $\lambda_{\operatorname{pos}}$ and $\lambda_{\operatorname{scale}}$ are set to 0.01 and 1 respectively.
@@ 4.4 Implementation Details
The FLAME meshes are initially obtained by VHAP tracker [47] from monocular videos. The animatable Gaussians are optimized with Adam [26] for 6,000 iterations, with learning rates of ${5}\mathrm{e}{-5}$,${1.7}\mathrm{e}{-2}$, ${1}\mathrm{e}{-3}$, ${2.5}\mathrm{e}{-3}$, and ${5}\mathrm{e}{-2}$ for splat position, scaling factor, rotation quaternion, color and opacity respectively. We perform adaptive densification if the position gradients are larger than $0.0002$ in every 300 iterations until 5,000 iterations are reached. We also remove Gaussians with opacity less than 0.005. During training, we also finetune the FLAME parameters using the learning rates ${1}\mathrm{e}{-6}$, ${1}\mathrm{e}{-5}$, and ${1}\mathrm{e}{-3}$ for translation, joint rotation, and expression coefficients. We randomly sample 4 novel viewpoints to calculate multi-view diffusion loss. The azimuth range is $[-90^{\circ},90^{\circ}]$, the elevation range is $[-30^{\circ},30^{\circ}]$, and the distance between the camera and the head avatar location is between 1.0 and 1.2 units.
@@
Our current implementation of avatar reconstruction takes about 12 hours and uses 32 GB of memory on a single A6000 GPU. After avatar reconstruction, our method takes 0.0026 seconds to render an image at 768$\times$768 resolution, i.e. 384.3fps.
@@ 5 Experiments
We compare our method against three state-of-the-art methods. GaussianAvatars[48] is designed to reconstruct head Gaussian splats rigged to FLAME model from multi-view videos captured in a studio environment. INSTA[91] employs a multi-resolution hash grid representation [42] for 3D head reconstruction. FlashAvatar [75] attaches a uniform 3D Gaussian field in the UV space of the FLAME mesh. Both INSTA and FlashAvatar are monocular-based methods.
@@ 5.1 Head Avatar Reconstruction
Ours 0.158 19.20 87.00 0.111 22.78 89.46
Table 1: Quantitative comparisons on dynamic avatar reconstruction and animation from monocular videos. Results are obtained by the average of twelve sequences of different subjects on the Nersemble dataset.
(a) (b) (c) (d) (e) (f)
Figure 4: Novel view synthesis from monocular videos on the NeRSemble dataset. (a) Input; (b) Ground Truth; (c) INSTA; (d) FlashAvatar; (e) GaussianAvatars; (f) Ours. Compared to state-of-the-art methods, our approach reconstructs unseen side facial regions in the inputs and consistently produces more favorable and consistent renderings from hold-out views.
(a) (b) (c) (d) (e) (f)
Figure 5: Head avatar reconstruction from monocular videos captured on commodity devices. (a) Input; (b) Ground truth of novel expressions; (c) INSTA; (d) FlashAvatar; (e) GaussianAvatars; (f) Ours. For each method, we display the fitting results of the input frame (top right) and novel view renderings of the input frame (bottom right). Given a front-facing sequence with limited head poses, all methods can accurately fit the observed frames. However, without effective priors to constrain unobserved regions, baseline methods struggle to generalize to novel views and poses.
@@
As shown in Fig. 4, front-facing monocular videos lack sufficient side-face information. Existing methods can only reconstruct observed regions and leave unobserved areas unconstrained. This often leads to artifacts in extreme hold-out views. In contrast, our approach leverages multi-view diffusion priors to constrain novel views, effectively completing missing regions and enhancing photorealism, while preserving identity and appearance consistency. In Tab. 1, our method surpasses all baselines across metrics, with notable gains of 1.69dB in PSNR and a 5.34% increase in SSIM for novel view synthesis.
@@
We also provide the comparisons on the Monocular Video dataset. Since single-view datasets lack ground-truth novel views, we evaluate animation results by applying expression parameters from hold-out frames. As shown in Tab. 2 , our method consistently surpasses others in all metrics. The qualitative results are presented in Fig. 5. While all methods fit well to the input view (top right), baseline methods struggle to generate plausible novel views (bottom right), especially for previously unseen face regions. Baseline methods show limited extrapolation capability, often resulting in artifacts or inconsistencies when animated to novel head poses.

### v2

@@ Abstract
We propose a novel approach for reconstructing animatable 3D Gaussian avatars from monocular videos captured by commodity devices like smartphones. Photorealistic 3D head avatar reconstruction from such recordings is challenging due to limited observations, which leaves unobserved regions under-constrained and can lead to artifacts in novel views. To address this problem, we introduce a multi-view head diffusion model, leveraging its priors to fill in missing regions and ensure view consistency in Gaussian splatting renderings. To enable precise viewpoint control, we use normal maps rendered from FLAME-based head reconstruction, which provides pixel-aligned inductive biases. We also condition the diffusion model on VAE features extracted from the input image to preserve facial identity and appearance details. For Gaussian avatar reconstruction, we distill multi-view diffusion priors by using iteratively denoised images as pseudo-ground truths, effectively mitigating over-saturation issues. To further improve photorealism, we apply latent upsampling priors to refine the denoised latent before decoding it into an image. We evaluate our method on the NeRSemble dataset, showing that GAF outperforms previous state-of-the-art methods in novel view synthesis. Furthermore, we demonstrate higher-fidelity avatar reconstructions from monocular videos captured on commodity devices. Project Page: https://tangjiapeng.github.io/projects/GAF
@@ 1 Introduction
To exploit multi-view diffusion priors for Gaussian head reconstruction, we employ a diffusion loss that uses iteratively denoised images as pseudo-ground truths for novel view renderings, instead of using a single-step score distillation sampling loss [58]. Compared to NeRF, 3D-GS enables real-time multi-view rendering, significantly reducing optimization time. Moreover, to improve the fidelity of the Gaussian renderings, we introduce a latent upsampler model to enhance facial details within the denoised latent before decoding it back to image space. As demonstrated in Fig. 1, our approach generates high-quality, photorealistic novel views of head avatars using only short monocular videos captured by a smartphone. Extensive comparisons with state-of-the-art methods show that our approach delivers higher fidelity and more view-consistent avatar reconstructions from monocular videos. Our GAF significantly outperforms state-of-the-art methods in novel view and expression synthesis on the NeRSemble dataset and other monocular datasets captured by commodity devices. Our contributions can be summarized as follows:
@@
  • We present a strategy to enhance the photorealism and cross-view consistency of Gaussian splatting by integrating latent upsampling and multi-view diffusion priors.
@@ 2 Related Work
To create photo-realistic animatable head avatars, several works [21, 110, 82, 110, 106, 92, 69, 60, 93] have combined NeRF/GS with 3D morphable models (3DMM). Our work is closely related to animatable Gaussian splats [60, 69], which attached splats to the triangles of FLAME mesh, and updated their properties by triangle deformations controlled by FLAME parameters. Although promising results have been achieved, they typically require multi-view setups with high-quality cameras in studio environments [46, 36]. Some works reconstruct avatars from monocular videos [106, 4, 110, 92]. However, they only reconstruct the visible region from inputs, lacking effective priors to complete missing areas. P4D [17], P4D-v2 [18], GAGAvatar [14], and MGGTalk [25] use feed-forward networks to predict expression-driven 3D heads represented by Gaussians or triplane features [56]. While achieving impressive head reenactments, they could exhibit obvious artifacts in novel views. To address this, we utilize multi-view diffusion priors to complete unobserved regions of the face, ensuring photorealistic renderings from extremely novel viewpoints.
@@
Diffusion models [73, 74, 75, 29] have been widely applied to 3D asset generation [100, 80, 9, 85, 49, 107]. Some methods leverage large-scale pretrained text-to-image priors [64, 67, 63] for 3D synthesis [41, 78] using score distillation sampling (SDS) loss [58] and its variants [89, 98, 50]. Several studies [28, 10, 109] applied SDS loss for text-driven 3D head avatar generation. For single-view 3D reconstruction, approaches like RelFusion [50], Magic123 [59], and Dreambooth3D [62] adapt text-to-image diffusion priors to specific objects to preserve identity. However, achieving consistent novel views from text prompts alone remains challenging without leveraging the input image. To address this, we learn multi-view image diffusion priors conditioned on a single image. Rather than using a single-step SDS loss, we use iteratively denoised images as pseudo-ground truths for novel view supervision, similar to ReconFusion [91]. We also introduce pretrained latent upsampler diffusion priors to improve photo-realism of pseudo-ground truths.
@@
Instead of text-to-image diffusion priors, some works learn image-conditioned novel view diffusion priors, which leverage the identity and appearance details of the input image to generate consistent novel views. 3DiM [90] pioneered a diffusion model for pose-conditioned novel view generation. Zero-123 finetuned StableDiffusion [64] on the Objaverse [15, 16] dataset to improve generalization ability. Some works jointly diffused multiple views from a single image via dense 3D self-attention or epipolar attention layers in the denoiser network, including MVDiffusion [81], SyncDreamer [45], Zero-123++ [70], Wonder3D [47], MVDream [71], ImageDream [88], Human3Diffusion [96], IM3D [51], CAT3D [22], and VideoMV [111]. These works primarily target general objects, but our focus is on human heads. We train multi-view diffusion models on a multi-view head video dataset to obtain more head-specific priors. By using normal maps from FLAME reconstruction as the camera pose condition, we introduce pixel-aligned inductive bias to enable more precise viewpoint control, which is crucial for consistent head Gaussian supervision.
@@ 3.2 Gaussian Avatars
$\vspace{-2mm}\vec{r}^{\prime}=\vec{R}\vec{r}$ (1)
$\vspace{-1mm}\vec{\mu}^{\prime}=k\vec{R}\vec{\mu}+\vec{T}$ (2)
$\vspace{-2mm}\vec{s}^{\prime}=k\vec{s}$ (3)
The adaptive Gaussian densification process operates within local space, where newly added Gaussians inherit binding relations from their original counterparts.
@@ 4.1 Normal map-conditioned Multi-view Head Latent Diffusion
To address missing regions in monocular videos with limited head coverage, one solution is to distill pre-trained text-to-image diffusion models to regularize novel view renderings. Personalized techniques like Dreambooth [66] allow for identity preservation by customizing diffusion models for specific objects we wish to reconstruct. However, personalized text-to-image diffusion models are not designed to capture the distribution of novel views from a single input image. To overcome this, we propose a novel view diffusion model that generates identity-preserved and appearance-coherent novel views conditioned on an input image. By denoising multiple novel views simultaneously, our approach enhances cross-view consistency. The multi-view head diffusion model is illustrated in Fig. 3.
@@
Figure 3: Multi-view latent head diffusion models. Given multi-view noisy image latents, we concatenate them with VAE latents of normal maps rendered from FLAME tracking. These combined inputs are processed by a 2D U-Net denoiser with attention blocks. To maintain 3D consistency, 3D attention blocks apply cross-attention across all views, integrating face identity and appearance details from the input image into the denoised latents while exchanging information between noisy latents across views.
@@ 4.2 Gaussian Avatars Reconstruction with Multi-view Diffusion Priors
We now seek to utilize the multi-view diffusion priors for head Gaussian reconstruction. A commonly used strategy is Score Distillation Sampling (SDS) loss [58], which performs one-step denoising. Due to the stochastic nature of the denoising process introduced by random noise levels and seeds, it contains noisy gradients that disturb 3D optimization. Consequently, it often causes over-saturated appearance issues in synthesized 3D assets. Wu et al. [91] found that after iteratively denoising a noisy image for multiple steps, we can obtain a deterministic output.
@@
At each iteration, we randomly select the $i$-th input frame $\vec{I}_{i}$ and its FLAME mesh $\mathcal{M}_{i}$. By randomly sampling 4 viewpoints $\{\phi_{j}\}_{j=1}^{4}$, then we can generate 4 novel views from $\vec{I}_{i}^{\operatorname{view}}=\mathcal{R}(\mathcal{O}_{i},\{\phi_{j}\})$, and render normal maps $\mathbf{N}_{i}=\mathcal{R}(\mathcal{M},\{\phi_{j}\})$ from $\mathcal{M}_{i}$. We can employ normal map-guided multi-view diffusion priors to regularize $\vec{I}_{i}^{\operatorname{view}}$. Concretely, we encode $\vec{I}_{i}^{\operatorname{view}}$ into latent features $\vec{z}$, which is perturbed with noise $\epsilon_{t}$ to obtain noisy features $\vec{z}_{t}$. The noise intensity of $\epsilon_{t}$ is controlled by the diffusion time step $t\sim[0.02,0.98]$. We then iteratively perform multiple denoising steps of latent diffusion until the final clean latent $\vec{z}_{0}$ is obtained. To accelerate the generation speed, we adopt the DDIM sampling strategy [72] by running $t/k$ intermediate steps, which can reduce the denoising steps by $k=20$ times. Then $\vec{z}_{0}$ is decoded back to 4 images $\{\vec{\hat{I}}_{i}^{j}\}$ that are served as pseudo supervision.
@@
To further enhance the facial details in $\vec{\hat{I}}_{i}^{\operatorname{view}}$, we use a pre-trained latent upsampler diffusion model [2] to super-resolve the denoised 4-view latents $\vec{z}_{0}$, from a resolution of 32 $\times$ 32 to 64 $\times$ 64. This super-resolution step allows the pseudo-image ground truths $\vec{\hat{I}}_{i}^{j}$ to reach a final resolution of 512$\times$512. We use 10-step DDIM sampling in our latent upsampler inference.
@@ 4.3 Loss Functions
$\vspace{-2mm}\mathcal{L}_{\operatorname{img}}=\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{SSIM}+\lambda_{3}\mathcal{L}_{LPIPS}$ (6)
where $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.1$. We also introduce splat position and scale regularization terms to penalize abnormally distributed splats. $\lambda_{\operatorname{pos}}$ and $\lambda_{\operatorname{scale}}$ are set to 0.01 and 1 respectively.
@@ 4.4 Implementation Details
The FLAME meshes are initially obtained by VHAP tracker [1] from monocular videos. The animatable Gaussians are optimized with Adam [35] for 6,000 iterations, with learning rates of ${5}\mathrm{e}{-5}$,${1.7}\mathrm{e}{-2}$, ${1}\mathrm{e}{-3}$, ${2.5}\mathrm{e}{-3}$, and ${5}\mathrm{e}{-2}$ for splat position, scaling factor, rotation quaternion, color and opacity respectively. We perform adaptive densification if the position gradients are larger than $0.0002$ in every 300 iterations until 5,000 iterations are reached. We also remove Gaussians with opacity less than 0.005. During training, we also finetune the FLAME parameters using the learning rates ${1}\mathrm{e}{-6}$, ${1}\mathrm{e}{-5}$, and ${1}\mathrm{e}{-3}$ for translation, joint rotation, and expression coefficients. At each iteration, we randomly sample 4 out of 15 novel viewpoints from the NeRSemble dataset to calculate multi-view diffusion loss.
@@
Our current implementation of avatar reconstruction takes about 12 hours and uses 32 GB of memory on a single A6000 GPU. After avatar reconstruction, Our method takes 0.016 seconds to render an image at 802$\times$550 resolution, i.e. 62 FPS.
@@ 5 Experiments
We compare our method against state-of-the-art methods for photo-realistic head reconstruction. PanoHead [3] and SphereHead [38] use 3D-GAN inversion to reconstruct static heads. They cannot reconstruct animatable avatars. Thus, they cannot be evaluated in novel expression synthesis. INSTA[110], FlashAvatar [92], and GA[60] optimize photo-realistic head avatars represented by Instant NGP or Gaussian Splatting. They overfit avatars to monocular observations. P4D-v2 [18] and GAGAvatar [14] reconstruct animatable 3D avatars from single images, that can then be reenacted with novel expressions.
@@ 5.1 Head Avatar Reconstruction
P4D-v2 [18] 0.161 16.82 85.44 0.142 18.72 86.63
GAGAvatar [14] 0.183 21.20 80.72 0.177 21.32 80.44
Ours 0.125 20.88 88.91 0.087 24.12 90.66
Table 1: Quantitative comparisons on dynamic avatar reconstruction and animation from monocular videos. Results are obtained by the average of twelve sequences of different subjects on the Nersemble dataset.  Best and 2nd-best are highlighted.
Input GT PanoHead SphereHead INSTA FlashAvatar GA P4D-v2 GAGAvatar Ours
Figure 4: Novel view synthesis from monocular videos on the NeRSemble dataset. Compared to state-of-the-art methods, our approach reconstructs unseen side facial regions in the inputs, better preserves facial identities, and consistently produces more favorable renderings from hold-out views.
Input (a) (b) (c) (d) (e)
Figure 5: Head avatar reconstruction from monocular videos captured on commodity devices. (a) Ground truth of novel expressions; (b) GA; (c) P4D-v2; (d) GAGAvatar; (e) Ours. For each method, we display the fitting results of the input frame (top right) and novel view renderings of the input frame (bottom right). Given a front-facing sequence with limited head poses, all methods can accurately reconstruct the observed frames. However, without effective priors to constrain unobserved regions, GA struggles to generalize to novel views and poses. P4D-v2 and GAGAvatar exhibit limitations in capturing complicated facial expressions or fine-grained details such as wrinkles.
@@
As shown in Fig. 4, front-facing monocular videos lack sufficient side-face information. INSTA, FlashAvatar, and GA can only reconstruct observed regions and leave unobserved areas unconstrained. This often leads to artifacts in extreme hold-out views. PanoHead [3] and SphereHead [38] fail to generate plausible renderings in extreme novel views. P4D-v2 [18] exhibits blurriness and artifacts in the mouth and ears, while GAGAvatar [14] suffers from identity drift due to weak constraints in less observed regions. In contrast, our method leverages multi-view diffusion priors to improve novel view synthesis, effectively completing missing regions and enhancing photorealism while preserving identity and appearance consistency. As shown in Tab. 1, our approach outperforms all baselines in LPIPS and SSIM and ranks second in PSNR on the task of novel view synthesis. Interestingly, novel view constraints also benefit novel expression synethesis, resulting in superior performance across all metrics.
@@
We also provide the comparisons on the Monocular Video dataset. Since single-view datasets lack ground-truth novel views, we evaluate animation results by applying expression parameters from hold-out frames. As shown in Tab. 2 , our method consistently surpasses others in all metrics. The qualitative results are presented in Fig. 5. While all methods can well reconstruct the input view (top right), GA [60] struggles to generate plausible novel views (bottom right), particularly for unseen face regions. It is also prone to artifacts and inconsistencies in novel pose animations. P4D-v2 [18] and GAGAvatar [14] have limited capacity to represent exaggerated expressions or fine details such as wrinkles.